OpenStudy (anonymous):

Let f(x) = 9x − 7, g(x) = |x|, h(x) = sq. root of x and k(x) = 3/x. Find and simplify the indicated composite function. (f o k)(x)

OpenStudy (anonymous):

I got 9-3/x-7 but I dont think thats right, and when I simplify it to 2-3/x it tells me my answers wrong

OpenStudy (zale101):

f(x) = 9x − 7 k(x) = 3/x (f o k)(x) is another way of saying (f(k(x))) (f(k(x))) =(f(3/x)) replace 3/x with x in the function of f(x) 9(3/x)-7